OpenStudy (anonymous):

A boy buys x eggs at at (x-8) pence each and (x-2) loafs of bread at (x-3) pence each. If the total bill is £1.75, how many eggs does he buy???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its (x-8) + (x-2)(x-3) = 1.75

OpenStudy (anonymous):

x-8 +x^2-5x+6 = 1.75 solve for x

OpenStudy (alexwee123):

if its (x-3) pence each it should be (x-8) + (x-2)/(x-3) = 1.75

OpenStudy (anonymous):

im missing an x attached to the (x-8)...should be x(x-8)

OpenStudy (anonymous):

and dont listen to alexwee he's wrong.

OpenStudy (anonymous):

still getting the wrong answer.

OpenStudy (anonymous):

x(x-8) + (x-2)(x-3) = 1.75 x^2 -8x +x^2 -5x+6 = 1.75...solve

OpenStudy (alexwee123):

x(x-8)+(x-2)(x-2)=1.75

OpenStudy (alexwee123):

dang

OpenStudy (alexwee123):

x-3*^

OpenStudy (anonymous):

what roots did you guys get?

OpenStudy (anonymous):

I really would like to know the correct answer

OpenStudy (kropot72):

x(x-8) + (x-2)(x-3) = 175 2x^2 - 13x + 6 = 175 2x^2 - 13x -169 = 0 (2x + 13)(x - 13) = 0 Therefore x = 13 the answer is 13 eggs were bought.

OpenStudy (alexwee123):

is a euro equivalent to 100 pence?

OpenStudy (anonymous):

@alexwee123 no

OpenStudy (kropot72):

An English pound = 100 new pence

OpenStudy (anonymous):

@kropot72 what happened to the 1.75???

OpenStudy (alexwee123):

he converted it into pence?

OpenStudy (anonymous):

sick

OpenStudy (anonymous):

why did you think to do that?

OpenStudy (kropot72):

The pound was converted to pence, the reason being that all the pricing is in pence.

OpenStudy (anonymous):

That why my teacher urges me to READ THE QUESTION TWICE>